# Scalable learning of macroscopic stochastic dynamics

### v1

@@ Appendix B Theoretical Analysis
$\delta\hat{\mathcal{L}}$ $=\mathbb{E}_{\mathbf{x}_{t},\delta t}\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}\big[-\text{tr}(\boldsymbol{\Lambda}^{-1}\mathbf{H})+\tfrac{1}{\delta t}\boldsymbol{\delta}^{T}\mathbf{H}\boldsymbol{\delta}-2\boldsymbol{\delta}^{T}\boldsymbol{\Lambda}\mathbf{h}\big],$
$\delta^{2}\hat{\mathcal{L}}$ $=\mathbb{E}_{\mathbf{x}_{t},\delta t}\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}\big[\text{tr}(\boldsymbol{\Lambda}^{-1}\mathbf{H}\boldsymbol{\Lambda}^{-1}\mathbf{H})+2\delta t\mathbf{h}^{T}\boldsymbol{\Lambda}\mathbf{h}-4\boldsymbol{\delta}^{T}\mathbf{H}\mathbf{h}\big].$
If $\boldsymbol{\mu}^{\ast},\boldsymbol{\Lambda}^{\ast}$ are the minimizer of $\hat{\mathcal{L}}$, we must have $\delta\hat{\mathcal{L}}=0$ for all the admissible $\mathbf{h}$ and $\mathbf{H}$. Since
@@
$=\mathbb{E}_{\mathbf{x}_{t},\delta t}\Big[\text{tr}\big(\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}\big[\tfrac{1}{\delta t}\boldsymbol{\delta}\boldsymbol{\delta}^{T}-\boldsymbol{\Lambda}^{-1}\big]\mathbf{H}\big)\Big]+\mathbb{E}_{\mathbf{x}_{t},\delta t}\Big[\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}\big[-2\boldsymbol{\delta}^{T}\boldsymbol{\Lambda}\big]\mathbf{h}\Big],$
we have:
$\quad\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}\big[\tfrac{1}{\delta t}\boldsymbol{\delta}\boldsymbol{\delta}^{T}-(\boldsymbol{\Lambda}^{\ast})^{-1}\big]=0\quad\textit{a.e.},$
$\quad\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}\big[-2\boldsymbol{\delta}^{T}\boldsymbol{\Lambda}\big]=\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}\big[-2(\mathbf{z}_{t+\delta t}-\mathbf{z}_{t}-\boldsymbol{\mu}^{\ast}(\mathbf{z}_{t})\delta t)^{T}\big]\boldsymbol{\Lambda}^{\ast}(\mathbf{z}_{t})=0\quad\textit{a.e.},$
hence
$\boldsymbol{\mu}^{\ast}(\mathbf{z}_{t})$ $=\tfrac{1}{\delta t}\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}[\mathbf{z}_{t+\delta t}-\mathbf{z}_{t}]\quad\textit{a.e.},$
$\boldsymbol{\Sigma}^{\ast}(\mathbf{z}_{t})$ $=(\boldsymbol{\Lambda}^{\ast})^{-1}(\mathbf{z}_{t})=\tfrac{1}{\delta t}\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}[\boldsymbol{\delta}\boldsymbol{\delta}^{T}]\quad\textit{a.e.}.$
@@
$\delta^{2}\hat{\mathcal{L}}$ $=\mathbb{E}_{\mathbf{x}_{t},\delta t}\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}\big[\text{tr}(\boldsymbol{\Lambda}^{-1}\mathbf{H}\boldsymbol{\Lambda}^{-1}\mathbf{H})+2\delta t\mathbf{h}^{T}\boldsymbol{\Lambda}\mathbf{h}\big]-4\mathbb{E}_{\mathbf{x}_{t},\delta t}\Big[\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}\big[(\mathbf{z}_{t+\delta t}-\mathbf{z}_{t}-\boldsymbol{\mu}^{\ast}(\mathbf{z}_{t})\delta t)^{T}\big]\mathbf{H}\mathbf{h}\Big]$
@@
$\delta^{2}\hat{\mathcal{L}}>0$ if either $\mathbf{H}$ or $\mathbf{h}$ is not equal to zero almost everywhere. Then $\boldsymbol{\mu}^{\ast}$ and $\boldsymbol{\Lambda}^{\ast}$ are the unique minimizer of functional $\hat{\mathcal{L}}$. Equivalently, $\boldsymbol{\mu}^{\ast}$ and $\boldsymbol{\Sigma}^{\ast}$ are the unique minimizer of $\mathcal{L}$. Similarly, we can compute the first-order variation and second-order variation of $\hat{\mathcal{L}}_{p}$:
@@
$\delta\hat{\mathcal{L}}_{p}$ $=\mathbb{E}_{\mathbf{x}_{t},\delta t}\mathbb{E}_{\mathcal{I}}\mathbb{E}_{\mathbf{x}_{t+\delta t,\mathcal{I}}|\mathbf{x}_{t}}\big[-\text{tr}(\boldsymbol{\Lambda}^{-1}\mathbf{H})+\tfrac{1}{K\delta t}\boldsymbol{\delta}_{\mathcal{I}}^{T}\mathbf{H}\boldsymbol{\delta}_{\mathcal{I}}-\tfrac{2}{K}\boldsymbol{\delta}_{\mathcal{I}}^{T}\boldsymbol{\Lambda}\mathbf{h}\big],$
$\delta^{2}\hat{\mathcal{L}}_{p}$ $=\mathbb{E}_{\mathbf{x}_{t},\delta t}\mathbb{E}_{\mathcal{I}}\mathbb{E}_{\mathbf{x}_{t+\delta t,\mathcal{I}}|\mathbf{x}_{t}}\big[\text{tr}(\boldsymbol{\Lambda}^{-1}\mathbf{H}\boldsymbol{\Lambda}^{-1}\mathbf{H})+\tfrac{2\delta t}{K}\mathbf{h}^{T}\boldsymbol{\Lambda}\mathbf{h}-\tfrac{4}{K}\boldsymbol{\delta}_{\mathcal{I}}^{T}\mathbf{H}\mathbf{h}\big].$
If $\boldsymbol{\mu}^{\dagger},\boldsymbol{\Lambda}^{\dagger}$ are the minimizer of $\hat{\mathcal{L}}_{p}$, we must have $\delta\hat{\mathcal{L}}_{p}=0$ for all the admissible $\mathbf{h}$ and $\mathbf{H}$. Since
$\delta\hat{\mathcal{L}}_{p}$ $=\mathbb{E}_{\mathbf{x}_{t},\delta t}\mathbb{E}_{\mathcal{I}}\mathbb{E}_{\mathbf{x}_{t+\delta t,\mathcal{I}}|\mathbf{x}_{t}}\big[-\text{tr}(\boldsymbol{\Lambda}^{-1}\mathbf{H})+\tfrac{1}{K\delta t}\boldsymbol{\delta}_{\mathcal{I}}^{T}\mathbf{H}\boldsymbol{\delta}_{\mathcal{I}}-\tfrac{2}{K}\boldsymbol{\delta}_{\mathcal{I}}^{T}\boldsymbol{\Lambda}\mathbf{h}\big]$
$=\mathbb{E}_{\mathbf{x}_{t},\delta t}\Big[\text{tr}\big(\mathbb{E}_{\mathcal{I}}\mathbb{E}_{\mathbf{x}_{t+\delta t,\mathcal{I}}|\mathbf{x}_{t}}\big[\tfrac{1}{K\delta t}\boldsymbol{\delta}_{\mathcal{I}}\boldsymbol{\delta}_{\mathcal{I}}^{T}-\boldsymbol{\Lambda}^{-1}\big]\mathbf{H}\big)\Big]+\mathbb{E}_{\mathbf{x}_{t},\delta t}\Big[\mathbb{E}_{\mathcal{I}}\mathbb{E}_{\mathbf{x}_{t+\delta t,\mathcal{I}}|\mathbf{x}_{t}}\big[-\tfrac{2}{K}\boldsymbol{\delta}_{\mathcal{I}}^{T}\boldsymbol{\Lambda}\big]\mathbf{h}\Big],$
@@
$\quad\mathbb{E}_{\mathcal{I}}\mathbb{E}_{\mathbf{x}_{t+\delta t,\mathcal{I}}|\mathbf{x}_{t}}[-\tfrac{2}{K}\boldsymbol{\delta}_{\mathcal{I}}^{T}\boldsymbol{\Lambda}^{\dagger}]=\mathbb{E}_{\mathcal{I}}\mathbb{E}_{\mathbf{x}_{t+\delta t,\mathcal{I}}|\mathbf{x}_{t}}[-\tfrac{2}{K}(\mathbf{z}_{t+\delta t,\mathcal{I}}-\mathbf{z}_{t}-\boldsymbol{\mu}^{\dagger}(\mathbf{z}_{t})\delta t)^{T}]\boldsymbol{\Lambda}^{\dagger}(\mathbf{z}_{t})=0\quad\textit{a.e.}$
$\quad\mathbb{E}_{\mathcal{I}}\mathbb{E}_{\mathbf{x}_{t+\delta t,\mathcal{I}}|\mathbf{x}_{t}}[\tfrac{1}{K\delta t}\boldsymbol{\delta}_{\mathcal{I}}\boldsymbol{\delta}_{\mathcal{I}}^{T}-(\boldsymbol{\Lambda}^{\dagger})^{-1}]=0\quad\textit{a.e.}.$
@@
$\boldsymbol{\mu}^{\dagger}(\mathbf{z}_{t})$ $=\tfrac{1}{\delta t}\mathbb{E}_{\mathcal{I}}\mathbb{E}_{\mathbf{x}_{t+\delta t,\mathcal{I}}|\mathbf{x}_{t}}[\mathbf{z}_{t+\delta t,\mathcal{I}}-\mathbf{z}_{t}]\quad\textit{a.e.},$
$\boldsymbol{\Sigma}^{\dagger}(\mathbf{z}_{t})$ $=(\boldsymbol{\Lambda}^{\dagger})^{-1}(\mathbf{z}_{t})=\tfrac{1}{K\delta t}\mathbb{E}_{\mathcal{I}}\mathbb{E}_{\mathbf{x}_{t+\delta t,\mathcal{I}}|\mathbf{x}_{t}}[\boldsymbol{\delta}_{\mathcal{I}}\boldsymbol{\delta}_{\mathcal{I}}^{T}]\quad\textit{a.e.}.$
@@
$\delta^{2}\hat{\mathcal{L}}_{p}$ $=\mathbb{E}_{\mathbf{x}_{t},\delta t}\mathbb{E}_{\mathcal{I}}\mathbb{E}_{\mathbf{x}_{t+\delta t,\mathcal{I}}|\mathbf{x}_{t}}\big[\text{tr}(\boldsymbol{\Lambda}^{-1}\mathbf{H}\boldsymbol{\Lambda}^{-1}\mathbf{H})+\tfrac{2\delta t}{K}\mathbf{h}^{T}\boldsymbol{\Lambda}^{\dagger}\mathbf{h}\big]-\tfrac{4}{K}\mathbb{E}_{\mathbf{x}_{t},\delta t}\mathbb{E}_{\mathcal{I}}\mathbb{E}_{\mathbf{x}_{t+\delta t,\mathcal{I}}|\mathbf{x}_{t}}[\boldsymbol{\delta}^{T}_{\mathcal{I}}\mathbf{H}\mathbf{h}]$
@@
$=\mathbb{E}_{\mathbf{x}_{t},\delta t}\mathbb{E}_{\mathcal{I}}\mathbb{E}_{\mathbf{x}_{t+\delta t,\mathcal{I}}|\mathbf{x}_{t}}\big[\text{tr}(\boldsymbol{\Lambda}^{-1}\mathbf{H}\boldsymbol{\Lambda}^{-1}\mathbf{H})+\tfrac{2\delta t}{K}\mathbf{h}^{T}\boldsymbol{\Lambda}^{\dagger}\mathbf{h}\big],$
$\delta^{2}\hat{\mathcal{L}}_{p}>0$ if either $\mathbf{H}$ or $\mathbf{h}$ is not equal to zero almost everywhere. Then $\boldsymbol{\mu}^{\dagger}$ and $\boldsymbol{\Lambda}^{\dagger}$ is the unique minimizer of functional $\hat{\mathcal{L}}$. Equivalently, $\boldsymbol{\mu}^{\dagger}$ and $\boldsymbol{\Sigma}^{\dagger}$ are the unique minimizer of $\mathcal{L}_{p}$.
Next we will prove $\boldsymbol{\mu}^{\ast}=\boldsymbol{\mu}^{\dagger}$, $\boldsymbol{\Sigma}^{\ast}=\boldsymbol{\Sigma}^{\dagger}$:
$\boldsymbol{\mu}^{\ast}(\mathbf{z}_{t})$ $=\tfrac{1}{\delta t}\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}[\mathbf{z}_{t+\delta t}-\mathbf{z}_{t}]$

### v2

@@ 1. Derivation of Minimizers
$\delta\hat{\mathcal{L}}$ $=\mathbb{E}_{\mathbf{x}_{t},\delta t}\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}\big[-\text{tr}(\boldsymbol{\Lambda}^{-1}\mathbf{H})$
$+\tfrac{1}{\delta t}\boldsymbol{\delta}^{T}\mathbf{H}\boldsymbol{\delta}-2\boldsymbol{\delta}^{T}\boldsymbol{\Lambda}\mathbf{h}\big],$
$\delta^{2}\hat{\mathcal{L}}$ $=\mathbb{E}_{\mathbf{x}_{t},\delta t}\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}\big[\text{tr}(\boldsymbol{\Lambda}^{-1}\mathbf{H}\boldsymbol{\Lambda}^{-1}\mathbf{H})$
$+2\delta t\mathbf{h}^{T}\boldsymbol{\Lambda}\mathbf{h}-4\boldsymbol{\delta}^{T}\mathbf{H}\mathbf{h}\big].$
If $\boldsymbol{\mu}^{\ast},\boldsymbol{\Lambda}^{\ast}$ are the minimizer of $\hat{\mathcal{L}}$, we must have $\delta\hat{\mathcal{L}}=0$ for all the admissible $\mathbf{h}$ and $\mathbf{H}$. By a direct calculation:
@@
$=\mathbb{E}_{\mathbf{x}_{t},\delta t}\Big[\text{tr}\big(\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}\big[\tfrac{1}{\delta t}\boldsymbol{\delta}\boldsymbol{\delta}^{T}-\boldsymbol{\Lambda}^{-1}\big]\mathbf{H}\big)\Big]$
$\quad+\mathbb{E}_{\mathbf{x}_{t},\delta t}\Big[\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}\big[-2\boldsymbol{\delta}^{T}\boldsymbol{\Lambda}\big]\mathbf{h}\Big],$
The stationary condition implies:
$\quad\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}\big[\tfrac{1}{\delta t}\boldsymbol{\delta}^{\ast}(\boldsymbol{\delta}^{\ast})^{T}-(\boldsymbol{\Lambda}^{\ast})^{-1}\big]=0\quad\textit{a.e.},$
$\quad\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}\big[-2(\boldsymbol{\delta}^{\ast})^{T}\boldsymbol{\Lambda}\big]$
$=\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}\big[-2(\Delta\mathbf{z}_{t}-\boldsymbol{\mu}^{\ast}(\mathbf{z}_{t})\delta t)^{T}\big]\boldsymbol{\Lambda}^{\ast}(\mathbf{z}_{t})$
$=0\quad\textit{a.e.},$
where $\boldsymbol{\delta^{\ast}}(\mathbf{z}_{t},\mathbf{z}_{t+\delta t})\coloneqq\Delta\mathbf{z}_{t}-\boldsymbol{\mu}^{\ast}(\mathbf{z}_{t})\delta t$. Solving these yields the unique minimizers for $\mathcal{L}$:
$\boldsymbol{\mu}^{\ast}(\mathbf{z}_{t})$ $=\tfrac{1}{\delta t}\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}[\Delta\mathbf{z}_{t}]=\tfrac{1}{\delta t}\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}[\mathbf{z}_{t+\delta t}-\mathbf{z}_{t}]\quad\textit{a.e.},$
$\boldsymbol{\Sigma}^{\ast}(\mathbf{z}_{t})$ $=(\boldsymbol{\Lambda}^{\ast})^{-1}(\mathbf{z}_{t})=\tfrac{1}{\delta t}\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}[\boldsymbol{\delta}^{\ast}(\boldsymbol{\delta}^{\ast})^{T}]\quad\textit{a.e.}.$
@@
$\delta^{2}\hat{\mathcal{L}}$ $=\mathbb{E}_{\mathbf{x}_{t},\delta t}\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}\big[\text{tr}(\boldsymbol{\Lambda}^{-1}\mathbf{H}\boldsymbol{\Lambda}^{-1}\mathbf{H})+2\delta t\mathbf{h}^{T}\boldsymbol{\Lambda}\mathbf{h}\big]$
$\quad-4\mathbb{E}_{\mathbf{x}_{t},\delta t}\Big[\mathbb{E}_{\mathbf{x}_{t+\delta t}|\mathbf{x}_{t}}\big[(\Delta\mathbf{z}_{t}-\boldsymbol{\mu}^{\ast}(\mathbf{z}_{t})\delta t)^{T}\big]\mathbf{H}\mathbf{h}\Big]$
@@
$\delta^{2}\hat{\mathcal{L}}>0$ if either $\mathbf{H}$ or $\mathbf{h}$ is not equal to zero almost everywhere. Then $\boldsymbol{\mu}^{\ast}$ and $\boldsymbol{\Lambda}^{\ast}$ are the unique minimizer of functional $\hat{\mathcal{L}}$. Equivalently, $\boldsymbol{\mu}^{\ast}$ and $\boldsymbol{\Sigma}^{\ast}$ are the unique minimizer of $\mathcal{L}$.
Similarly, we can compute the first-order variation and second-order variation of $\hat{\mathcal{L}}_{p}$:
@@
$\delta\hat{\mathcal{L}}_{p}$ $=\mathbb{E}_{\mathbf{x}_{t},\delta t}\mathbb{E}_{\mathcal{I}}\mathbb{E}_{q(\mathbf{x}_{t+\delta t,\mathcal{I}}|\mathbf{x}_{t})}\big[-\text{tr}(\boldsymbol{\Lambda}^{-1}\mathbf{H})$
$\quad+\tfrac{1}{K\delta t}\boldsymbol{\delta}_{\mathcal{I}}^{T}\mathbf{H}\boldsymbol{\delta}_{\mathcal{I}}-\tfrac{2}{K}\boldsymbol{\delta}_{\mathcal{I}}^{T}\boldsymbol{\Lambda}\mathbf{h}\big],$
$\delta^{2}\hat{\mathcal{L}}_{p}$ $=\mathbb{E}_{\mathbf{x}_{t},\delta t}\mathbb{E}_{\mathcal{I}}\mathbb{E}_{q(\mathbf{x}_{t+\delta t,\mathcal{I}}|\mathbf{x}_{t})}\big[\text{tr}(\boldsymbol{\Lambda}^{-1}\mathbf{H}\boldsymbol{\Lambda}^{-1}\mathbf{H})$
$\quad+\tfrac{2\delta t}{K}\mathbf{h}^{T}\boldsymbol{\Lambda}\mathbf{h}-\tfrac{4}{K}\boldsymbol{\delta}_{\mathcal{I}}^{T}\mathbf{H}\mathbf{h}\big].$
If $\boldsymbol{\mu}^{\dagger},\boldsymbol{\Lambda}^{\dagger}$ are the minimizer of $\hat{\mathcal{L}}_{p}$, we must have $\delta\hat{\mathcal{L}}_{p}=0$ for all the admissible $\mathbf{h}$ and $\mathbf{H}$. Let $\boldsymbol{\delta}_{\mathcal{I}}^{\dagger}(\mathbf{z}_{t},\mathbf{z}_{t+\delta,\mathcal{I}})\coloneqq\Delta\mathbf{z}_{t,\mathcal{I}}-\boldsymbol{\mu}^{\dagger}(\mathbf{z}_{t})\delta t$, since
$\delta\hat{\mathcal{L}}_{p}$ $=\mathbb{E}_{\mathbf{x}_{t},\delta t}\mathbb{E}_{\mathcal{I}}\mathbb{E}_{q(\mathbf{x}_{t+\delta t,\mathcal{I}}|\mathbf{x}_{t})}\big[-\text{tr}(\boldsymbol{\Lambda}^{-1}\mathbf{H})$
$\quad+\tfrac{1}{K\delta t}\boldsymbol{\delta}_{\mathcal{I}}^{T}\mathbf{H}\boldsymbol{\delta}_{\mathcal{I}}-\tfrac{2}{K}\boldsymbol{\delta}_{\mathcal{I}}^{T}\boldsymbol{\Lambda}\mathbf{h}\big]$
$=\mathbb{E}_{\mathbf{x}_{t},\delta t}\Big[\text{tr}\big(\mathbb{E}_{\mathcal{I}}\mathbb{E}_{q(\mathbf{x}_{t+\delta t,\mathcal{I}}|\mathbf{x}_{t})}\big[\tfrac{1}{K\delta t}\boldsymbol{\delta}_{\mathcal{I}}\boldsymbol{\delta}_{\mathcal{I}}^{T}-\boldsymbol{\Lambda}^{-1}\big]\mathbf{H}\big)\Big]$
$\quad+\mathbb{E}_{\mathbf{x}_{t},\delta t}\Big[\mathbb{E}_{\mathcal{I}}\mathbb{E}_{q(\mathbf{x}_{t+\delta t,\mathcal{I}}|\mathbf{x}_{t})}\big[-\tfrac{2}{K}\boldsymbol{\delta}_{\mathcal{I}}^{T}\boldsymbol{\Lambda}\big]\mathbf{h}\Big],$
@@
$\quad\mathbb{E}_{\mathcal{I}}\mathbb{E}_{q(\mathbf{x}_{t+\delta t,\mathcal{I}}|\mathbf{x}_{t})}[-\tfrac{2}{K}(\boldsymbol{\delta}_{\mathcal{I}}^{\dagger})^{T}\boldsymbol{\Lambda}^{\dagger}]$
$=\mathbb{E}_{\mathcal{I}}\mathbb{E}_{q(\mathbf{x}_{t+\delta t,\mathcal{I}}|\mathbf{x}_{t})}[-\tfrac{2}{K}(\Delta\mathbf{z}_{t,\mathcal{I}}-\boldsymbol{\mu}^{\dagger}(\mathbf{z}_{t})\delta t)^{T}]\boldsymbol{\Lambda}^{\dagger}(\mathbf{z}_{t})$
$=0\quad\textit{a.e.}$
$\quad\mathbb{E}_{\mathcal{I}}\mathbb{E}_{q(\mathbf{x}_{t+\delta t,\mathcal{I}}|\mathbf{x}_{t})}[\tfrac{1}{K\delta t}\boldsymbol{\delta}_{\mathcal{I}}^{\dagger}(\boldsymbol{\delta}_{\mathcal{I}}^{\dagger})^{T}-(\boldsymbol{\Lambda}^{\dagger})^{-1}]=0\quad\textit{a.e.}.$
@@
$\boldsymbol{\mu}^{\dagger}(\mathbf{z}_{t})$ $=\tfrac{1}{\delta t}\mathbb{E}_{\mathcal{I}}\mathbb{E}_{q(\mathbf{x}_{t+\delta t,\mathcal{I}}|\mathbf{x}_{t})}[\Delta\mathbf{z}_{t,\mathcal{I}}]$
$=\tfrac{1}{\delta t}\mathbb{E}_{\mathcal{I}}\mathbb{E}_{q(\mathbf{x}_{t+\delta t,\mathcal{I}}|\mathbf{x}_{t})}[\mathbf{z}_{t+\delta t,\mathcal{I}}-\mathbf{z}_{t}]\quad\textit{a.e.},$
$\boldsymbol{\Sigma}^{\dagger}(\mathbf{z}_{t})$ $=(\boldsymbol{\Lambda}^{\dagger})^{-1}(\mathbf{z}_{t})=\tfrac{1}{K\delta t}\mathbb{E}_{\mathcal{I}}\mathbb{E}_{q(\mathbf{x}_{t+\delta t,\mathcal{I}}|\mathbf{x}_{t})}[\boldsymbol{\delta}_{\mathcal{I}}^{\dagger}(\boldsymbol{\delta}_{\mathcal{I}}^{\dagger})^{T}]\quad\textit{a.e.}.$
@@
$\delta^{2}\hat{\mathcal{L}}_{p}$ $=\mathbb{E}_{\mathbf{x}_{t},\delta t}\mathbb{E}_{\mathcal{I}}\mathbb{E}_{q(\mathbf{x}_{t+\delta t,\mathcal{I}}|\mathbf{x}_{t})}\big[\text{tr}(\boldsymbol{\Lambda}^{-1}\mathbf{H}\boldsymbol{\Lambda}^{-1}\mathbf{H})+\tfrac{2\delta t}{K}\mathbf{h}^{T}\boldsymbol{\Lambda}^{\dagger}\mathbf{h}\big]$
$\quad-\tfrac{4}{K}\mathbb{E}_{\mathbf{x}_{t},\delta t}\mathbb{E}_{\mathcal{I}}\mathbb{E}_{q(\mathbf{x}_{t+\delta t,\mathcal{I}}|\mathbf{x}_{t})}[\boldsymbol{\delta}^{T}_{\mathcal{I}}\mathbf{H}\mathbf{h}]$
@@
$=\mathbb{E}_{\mathbf{x}_{t},\delta t}\mathbb{E}_{\mathcal{I}}\mathbb{E}_{q(\mathbf{x}_{t+\delta t,\mathcal{I}}|\mathbf{x}_{t})}\big[\text{tr}(\boldsymbol{\Lambda}^{-1}\mathbf{H}\boldsymbol{\Lambda}^{-1}\mathbf{H})+\tfrac{2\delta t}{K}\mathbf{h}^{T}\boldsymbol{\Lambda}^{\dagger}\mathbf{h}\big],$
2. Error Analysis
We now bound the difference between the estimators. To bound $\|\boldsymbol{\mu}^{\ast}-\boldsymbol{\mu}^{\dagger}\|_{\infty}$, we use the property of Total Variation distance. For any random variable $X$ bounded by $B$, we have
[…]
Next, since the encoder $\boldsymbol{\varphi}$ is uniformly bounded by $M$, it follows that $\|\Delta\mathbf{z}_{t,\mathcal{I}}\|_{\infty}\leq 2M$. We therefore obtain:
$\quad\lVert\boldsymbol{\mu}^{\dagger}(\mathbf{z}_{t})-\boldsymbol{\mu}^{\ast}(\mathbf{z}_{t})\rVert$